OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (triciaal):

nth term = a + (n-1)d where a is the 1st term and d is the common difference

OpenStudy (anonymous):

I don't know what the first term is

OpenStudy (triciaal):

is this all of the question? You have a14 and a15 seems like these are not the terms either

OpenStudy (anonymous):

That is all of the question @triciaal

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 a an = -579 + 42(n + 1) b an = -579 + 42(n - 1) c an = -579 - 42(n + 1) d an = -579 - 42(n - 1)

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

@satellite73

OpenStudy (triciaal):